OpenStudy (anonymous):

what is the proportion? 100 feet is to 140 feet as 15 feet is to 21 feet

OpenStudy (anonymous):

5/7

OpenStudy (anonymous):

BTW, you get that by simplifying 100/140

OpenStudy (anonymous):

I think you are looking for 5/7. ie, (100/20)/(140/20)=5/7 and (15/3)/(21/3)=5/7